# Glimpse: Generalized Local Imaging with MLPs

### v1

@@ III-D Resolution-agnostic Memory Usage in Training
To simplify notation, we denote the entire Glimpse pipeline described above by $\hat{{\mathbf{f}}}=\text{{Glimpse}}_{\phi}({\mathbf{x}},{\mathbf{s}})$. The inputs are the target pixel coordinates ${\mathbf{x}}=(x,y)$ and the sinogram ${\mathbf{s}}$; the output is an estimate of ${\mathbf{f}}(x,y)$. The parameters $\phi$ denote the trainable parameters of Glimpse including the MLP weights $\theta$, the projection angles $\{\alpha_{m}\}_{m=1}^{r}$, the adaptive filter ${\mathbf{h}}$ and the patch receptive field scale $d$. We consider a set of training data $\{({\mathbf{s}}_{i},{\mathbf{f}}_{i})\}_{i=1}^{L}$ from the noisy sinograms and images. We optimize the Glimpse parameters $\phi$ using gradient-based optimization to minimize
$\phi^{*}=\operatorname*{argmin}_{\phi}\sum_{i=1}^{N^{2}}\sum_{j=1}^{L}|\text{{Glimpse}}_{\phi}({\mathbf{x}}_{i},{\mathbf{s}}_{j})-{\mathbf{f}}_{j}({\mathbf{x}}_{i})|^{2}.$ (10)
@@
At inference time, we simply evaluate the image intensity at any pixel as $\hat{{\mathbf{f}}}_{\text{test}}({\mathbf{x}})=\text{{Glimpse}}_{\phi^{*}}({\mathbf{x}},{\mathbf{s}}_{\text{test}})$. One major advantage of Glimpse compared to CNNs like the U-Net and LPD is its memory and compute complexity. CNN-based models exhibit memory requirements that scale directly with image resolution, making them prohibitively expensive for realistic image resolutions. As shown in (10), Glimpse can be trained using stochastic gradient-based optimizers with the flexibility to select mini-batches from both the objects and pixels. This adaptability in mini-batch pixel selection grants a memory footprint agnostic to resolution making Glimpse suitable for training on realistic image resolutions like $1024\times 1024$ and higher.
@@ IV Experiments
We simulate parallel-beam X-ray CT with $r=30$ projections uniformly distributed around the object. We consider 30dB measurement noise. The reconstruction quality is quantified using the peak signal-to-noise ratio (PSNR) and Structural Similarity Index (SSIM) [38]. We compare the performance of Glimpse with successful CNN-based models: the U-Net [2], learned gradient scheme (LGS) [7] and learned primal-dual (LPD) [8] for sparse view CT image reconstruction. We use 35820 training samples of chest images from the LoDoPaB-CT dataset [39] in resolution $128\times 128$. Model performance is assessed on 64 in-distribution test samples of chest images, while 16 OOD brain images [40] are included to evaluate the generalization capability of the models. For further information regarding the network architectures and training details please refer to Section S-1.
In Section IV-A, we conduct a comparative analysis of Glimpse against CNN-based models for sparse view CT image reconstruction on both in-distribution and OOD data. In Section IV-B, we analyze the computational efficiency of the aforementioned models. Section IV-C addresses the uncalibrated and blind scenarios. We analyze the learned filters ${\mathbf{h}}$ across different measurement noise levels in Section IV-D. Finally in Section IV-E, we study performance of Glimpse for different patch sizes.
@@ IV-A Sparse view CT Image Reconstruction
TABLE I: Comparative analysis of different models for sparse view CT image reconstruction
@@
(b) Memory usage and training time
@@
The upper row of Figure 4 and Table I(a) show the performance of different models on in-distribution test samples of chest images. This experiment shows that Glimpse, by leveraging only a single MLP network, can outperform successful CNNs like U-Net and achieve comparable performance with state-of-the-art LGS and LPD methods.
The lower row of Figure 4 and Table I(a) provide a comparative analysis of the performance of various models trained on chest images when applied to OOD brain images. This experiment demonstrates that while U-Net excels on in-distribution samples, its performance significantly deteriorates on OOD data. This decline in performance may be attributed to the U-Net’s large receptive field which results in overfitting to large-scale features. We emphasize that this large receptive field is important to capture information delocalized in the filtered backprojection image and achieve strong performance. Indeed, reducing the receptive field, while a potential remedy for poor generalization on OOD data, may compromise the reconstruction quality even on in-distribution test data.
On the contrary, Glimpse demonstrates remarkable performance on OOD data. Although LPD’s performance on OOD data is sometimes comparable or slightly better than that of Glimpse, it comes at a very high memory and compute cost due to the repeated application of the forward operator and its adjoint in the network architecture.
@@ IV-B Computational Efficiency
The fact that LPD far outperforms the U-Net on OOD data is a testament to the benefits of incorporating the forward operator in the architecture. On the other hand, as evident from Table I(b), it comes at a cost of unfavorable training time and memory footprint which rapidly worsens with resolution. Table I(b) shows that CNN-based models may become impractical already at resolutions like $512\times 512$, even on GPUs with 80GB memory.
On the other hand, Glimpse is computationally efficient; the memory usage remains almost unaffected by image resolution. Remarkably, Glimpse can be trained with only 5GB memory in less than a day, even when dealing with resolutions of $1024\times 1024$ and higher. Figure 5 shows the performance of Glimpse on in-distribution and OOD samples in resolution $512\times 512$ where we considered 40dB measurement noise. This experiment demonstrates that even a small MLP can achieve reconstructions with near state-of-the-art quality in realistic high resolutions while maintaining a rather modest memory footprint.
@@
IV-C Uncalibrated Imaging and Blind Inversion
@@
In this section, we assess the performance of Glimpse in situations with calibration mismatch. In the following experiments, we place $r=30$ sensors uniformly around the object at angles $\alpha=0^{\circ},6^{\circ},...,174^{\circ}$. We conduct a comparative analysis of three models: 1) Glimpse (vanilla), featuring no learnable sensor geometry, 2) Glimpse (LSG), incorporating the proposed learned sensor geometry outlined in Section III-B, and 3) Glimpse (calibrated), operating under ideal conditions with no model mismatch (informed with correct projection angles). We let the Glimpse (LSG) learn the projection angles from the training data where the optimized values $\{\alpha_{m}\}_{i=1}^{r}$ obtained through training can provide a reliable estimate of the actual projection angles.
Uncalibrated system with random sensor shifts: As shown in Figure 5(a), we randomly perturb projection angles by a normally distributed error so that $\alpha_{i}^{\text{given}}=\mathcal{N}(\alpha_{i},\sigma^{2})$; we set $\sigma=2^{\circ}$. Figure 5(b) shows the estimated angles obtained through training—Glimpse (LSG) accurately recovers the angles with 30dB measurement noise. As shown in Figure 5(c), this accurate estimation of projection angles results in high-quality reconstructions by Glimpse (LSG) comparable with the network trained in an ideal calibrated system.
(a) Given projection angles
(b) Estimated projection angles
(c) Reconstructions
Figure 7: Estimated sensor geometry by Glimpse (LSG) and reconstructions for an uncalibrated system with a fixed $3^{\circ}$ sensor shift; as expected, the learnable sensor geometry can effectively learn the projection angles and exhibits excellent robustness.
Uncalibrated system with fixed sensor shifts: As depicted in Figure 6(a), we consider an uncalibrated system where the provided information about the projection angles deviates by $3^{\circ}$ from the actual positions, namely, $\alpha_{i}^{\text{given}}=\alpha_{i}+3^{\circ}$. As depicted in Figure 6(b), the estimated projection angles by the Glimpse (LSG) showcase its capability to accurately recover sensor positions from the training data. Figure 6(c) demonstrates the reconstructions for in-distribution and OOD data. This experiment shows that the performance of Glimpse (vanilla) sharply dropped in this uncalibrated system, particularly on OOD data. While its MLP module attempts to compensate for the model mismatch and succeeds to some extent on in-distribution chest images, this compensation comes at the cost of overfitting the training data, resulting in poor generalization to OOD samples. This experiment showcases Glimpse (LSG) as a robust deep learning model under model mismatch, well-suited for handling real-world challenges.
@@
Blind inversion with no information from projection angles: We consider the blind scenario where the model operates without any prior knowledge of the sensor geometry making inversion challenging. As shown in Figure 7(a), we inevitably initialize the projection angles $\{\alpha_{m}\}_{i=1}^{r}$ in the Glimpse (LSG) architecture with random values.
The estimated projection angles are shown in Figure 7(b), highlighting Glimpse (LSG)’s impressive ability for data-driven sensor geometry estimation. Figure 7(c) presents the reconstructions achieved by Glimpse in both its vanilla and LSG versions. As expected, FBP and the Glimpse (vanilla) show poor reconstructions due to the missing sensor geometry information. On the other hand, Glimpse (LSG) could accurately reconstruct both in-distribution and OOD samples. Remarkably, these results are comparable to those achieved by the calibrated Glimpse with informed projection angles. Table II provides quantitative results averaged on 64 test samples.
TABLE II: Model performance in PSNR (dB) averaged on 64 test samples for models with different levels of sensor geometry information: 1) calibrated, complete information 2) uncalibrated, with fixed and random sensor shifts 3) blind, no information.
(a) In-distribution chest samples
FBP Glimpse (vanilla) Glimpse (LSG)
Calibrated (complete info) 17.0 30.9 30.9
Uncalibrated (fixed shifts) 15.9 30.1 30.9
Uncalibrated (random shifts) 16.9 30.6 30.9
Blind (No info) 12.4 26.7 30.5
(b) OOD brain samples
FBP Glimpse (vanilla) Glimpse (LSG)
Calibrated (complete info) 17.1 25.1 25.1
Uncalibrated (fixed shifts) 15.6 20.0 25.1
Uncalibrated (random shifts) 16.1 23.2 25.0
Blind (No info) 10.8 14.9 24.1
IV-D Learned Filter
Figure 9: The learned filter for datasets with different noise levels, all the filtered are initialized by Ram-Lak filter in Glimpse architecture. By increasing the noise level, the filter assigns smaller amplitudes for high-frequencies to suppress the noise and aligns with the optimality of the Ram-Lak filter for noise-free complete measurements shown in Section S-2.
In this section, we analyze the behavior of the learned filter obtained through training of Glimpse introduced in section III-C across datasets with different measurement noise levels. This analysis provides useful signal processing insights into how the properties of the learned filter are influenced by varying noise levels. The frequency response of the learned filters is depicted in Figure 9, alongside well-known hand-crafted filters such as Ram-Lak, Shepp-Logan, and hamming filters. These learned filters are derived from Glimpse training on datasets characterized by different noise levels.
As expected and discussed in Section S-2, the learned filter for noise-free measurements exhibits behaviors akin to the Ram-Lak filter, displaying high values in high frequencies. Intriguingly, as the noise level increases (by decreasing the noise SNR), the filter progressively takes smaller values in high frequencies to suppress the noise. This experiment demonstrates that Glimpse is capable of autonomously adapting the characteristics of the filter according to the noise level observed in the training data. Reconstructed images for different noise levels are presented in Figure 10.
(a) In-distribution chest samples
(b) Out-of-distribution brain samples
Figure 10: Glimpse performance on in-distribution and OOD data for different measurement noise levels.
IV-E Ablation Study on Patch Size
In this section, we investigate the impact of varying the number of neighboring pixels $K=C^{2}$, on the Glimpse’s performance. Table III shows the performance of Glimpse trained with different numbers of neighboring pixels $K$ on in-distribution and OOD samples. The table demonstrates that the reconstruction quality tends to reach a saturation point beyond a certain patch size. This observation provides insights into understanding the upper limit of local information crucial for accurate reconstructions.
TABLE III: Reconstruction quality in PSNR (dB) for Glimpse trained with various window sizes
Patch size ($K=C^{2}$) In-distribution OOD Num params
$1\times 1$ 25.6 18.3 280k
$3\times 3$ 30.2 24.2 345k
$5\times 5$ 30.7 24.9 470k
$7\times 7$ 30.8 24.9 650k
$9\times 9$ 30.9 25.1 900k
$11\times 1$ 11 30.9 25.0 1200k
V Limitations and Conclusion
In this paper, we use the natural notion of locality for CT image reconstruction, moving away from CNNs that reconstruct the image as a whole. Our approach adopts a coordinate-based strategy, focusing on processing the sinusoidal curves in the sinogram associated with each pixel using a small MLP. Our results demonstrate that this localized processing framework exhibits robust performance on OOD data while maintaining nearly constant memory requirements across different resolutions; thereby being computationally efficient even at realistic high resolutions. Furthermore, our fully differentiable architecture presents versatility in addressing real-world challenges in medical imaging. Specifically, we showed its applicability in learning sensor geometry for uncalibrated systems and adapting the learned filter according to the noise level observed in training data.
The proposed model comes with certain limitations. While the memory required by Glimpse hardly varies with image resolution, memory and computing costs increase as the number of projections $r$ increases with the current MLP network. A promising alternative to the standard MLP architecture is the use of a mixture-of-experts layers [41, 42, 43], which selectively employ smaller MLPs for processing inputs. This mixture-of-experts approach is an effective drop-in replacement for standard MLP layers of language transformers [44] and vision transformers [26], as they increase the capacity of the layers without increasing the computational cost proportionally.
Moreover, while specifically designed for CT imaging, our localized architecture can integrate with various imaging problems involving line integrals in forward operators like photoacoustic [45, 46] and cryo-electron tomography (cryoET) [47]. In addition, although we introduced Glimpse specifically for 2D images, its adaptation to 3D shapes is a straightforward extension that can address the fundamental challenges for medical applications. This extension is particularly interesting given the ability of Glimpse to operate locally and its near-fixed memory requirement across resolution, which makes it an ideal choice for large 3D objects.
Supplementary Information
S-1 Network Architecture and Training Details
S-2 Optimal Filter for FBP Reconstruction

### v2

@@ III-C Resolution-agnostic Memory Usage in Training
To simplify notation, we denote the entire Glimpse pipeline described above by $\hat{{\mathbf{f}}}({\mathbf{x}})=\text{{Glimpse}}_{\phi}({\mathbf{x}},{\mathbf{s}})$. The inputs are the target pixel coordinates ${\mathbf{x}}=(x,y)$ and the sinogram ${\mathbf{s}}$; the output is an estimate of ${\mathbf{f}}(x,y)$. The parameters $\phi$ denote the trainable parameters of Glimpse including the MLP weights $\theta$, the projection angles $\{\alpha_{m}\}_{m=1}^{r}$ (see Appendix -B), the adaptive filter ${\mathbf{h}}$ and the window receptive field scale $d$. We consider a set of training data $\{({\mathbf{s}}_{i},{\mathbf{f}}_{i})\}_{i=1}^{L}$ from the noisy sinograms and images. We optimize the Glimpse parameters $\phi$ using gradient-based optimization to minimize
$\phi^{*}=\operatorname*{argmin}_{\phi}~{}\sum_{i=1}^{N^{2}}\sum_{j=1}^{L}|\text{{Glimpse}}_{\phi}({\mathbf{x}}_{i},{\mathbf{s}}_{j})-{\mathbf{f}}_{j}({\mathbf{x}}_{i})|^{2}.$ (10)
@@
At inference time, we simply evaluate the image intensity at any pixel as $\hat{{\mathbf{f}}}_{\text{test}}({\mathbf{x}})=\text{{Glimpse}}_{\phi^{*}}({\mathbf{x}},{\mathbf{s}}_{\text{test}})$. One major advantage of Glimpse compared to CNNs like U-Net and LPD is its memory and compute complexity. CNN-based models exhibit memory requirements that scale directly with image resolution, making them prohibitively expensive for realistic image resolutions. As shown in (10), Glimpse can be trained using stochastic gradient-based optimizers with the flexibility to select mini-batches from both the objects and pixels. This adaptability in mini-batch pixel selection grants a memory footprint agnostic to resolution making Glimpse suitable for training on realistic image resolutions like $1024\times 1024$ and higher.
@@ IV Experiments
We simulate parallel-beam X-ray CT with $r=30$ projections uniformly distributed around the object with additive Gaussian noise to reach a signal-to-noise ratio (SNR) of 30 dB. The reconstruction quality is quantified using the peak signal-to-noise ratio (PSNR) and Structural Similarity Index (SSIM) [40]. We compare the performance of Glimpse with successful CNN-based models: U-Net [2], learned gradient scheme (LGS) [7] and learned primal-dual (LPD) [8] for sparse view CT image reconstruction. We use 35820 training samples of chest images from the LoDoPaB-CT dataset [41] in resolution $128\times 128$. Model performance is assessed on 64 in-distribution test samples of chest images, while 16 OOD brain images [42] are included to evaluate the generalization capability of the models. For further information regarding the network architectures and training details please refer to Section -A.
In Section IV-A, we compare Glimpse to CNN-based models for sparse view CT image reconstruction on both in-distribution and OOD data. In Section IV-B, we analyze the computational efficiency of the aforementioned models. We analyze the learned filters ${\mathbf{h}}$ across different measurement noise levels in Section IV-C. In Appendix -B we consider the uncalibrated and blind scenarios.
@@ IV-A Sparse view CT Image Reconstruction
TABLE I: Comparison of different models for sparse view CT image reconstruction
@@
(b) Memory usage and training time (batch size 64)
@@
The upper row of Figure 4 and Table I(a) show the performance of different models on in-distribution test samples of chest images. This experiment shows that Glimpse, by leveraging only a single MLP network, can outperform successful CNNs like U-Net and achieve comparable performance with LGS and LPD methods.
The lower row of Figure 4 and Table I(a) shows a comparison of the performance of various models trained on chest images when applied to OOD brain images. This experiment demonstrates that while U-Net excels on in-distribution samples, its performance significantly deteriorates on OOD data.
On the contrary, Glimpse shows strong performance on OOD data. Although LPD’s performance on OOD data is sometimes comparable or slightly better than that of Glimpse, it comes at a very high memory and compute cost due to the repeated application of the forward operator and its adjoint in the network architecture; we analyze this in the next section.
@@ IV-B Computational Efficiency
The fact that LPD far outperforms U-Net on OOD data is a testament to the benefits of incorporating the forward operator in the architecture. On the other hand, as evident from Table I(b), it comes at the cost of unfavorable training time and memory footprint which rapidly worsens with resolution. Table I(b) shows that CNN-based models may become impractical already at resolutions like $512\times 512$, even on GPUs with 80GB memory.
On the other hand, Glimpse is computationally efficient; the memory usage remains almost unaffected by image resolution. Remarkably, Glimpse can be trained with only 5GB memory in less than a day, even when dealing with resolutions of $1024\times 1024$ and higher. Figure 5 shows the performance of Glimpse on in-distribution and OOD samples in resolution $512\times 512$ where we considered 40dB measurement noise. This experiment demonstrates that a relatively small MLP, with almost 10 times fewer parameters than a standard U-Net, can achieve strong performance in realistic high resolutions while maintaining a rather modest memory footprint.
@@
IV-C Learned Filter
Figure 6: The learned filter for datasets with different noise levels, all the filtered are initialized by Ram-Lak filter in Glimpse architecture. By increasing the noise level, the filter assigns smaller amplitudes for high-frequencies to suppress the noise and aligns with the optimality of the Ram-Lak filter for noise-free complete measurements shown in Section -C.
In this section, we analyze the behavior of the learned filter obtained through training of Glimpse introduced in section III-B across datasets with different measurement noise levels. This analysis provides useful signal processing insights into how the properties of the learned filter are influenced by varying noise levels.
In Figure 6 we show the frequency response of the learned filters, alongside with standard hand-crafted filters such as Ram-Lak, Shepp-Logan, and Hamming filters. These learned filters are derived from Glimpse training on datasets characterized by different levels of noise.
As expected from the discussion in Appendix -C, the learned filter for noise-free measurements is similar to the Ram–Lak filter, with a relatively high amplitude in high frequencies. As the noise level increases (by decreasing the noise SNR), the filter progressively takes smaller values in high frequencies to suppress the noise. This confirms that Glimpse can autonomously adapt the characteristics of the filter according to the noise level observed in the training data.
V Limitations and Conclusion
We used a natural notion of locality for CT which is adapted to sinogram geometry. This is different from CNNs that reconstruct the image as a whole, and where the notion of locality (at small scales) is in the sense of real image space. Our approach adopts a coordinate-based strategy, focusing on processing the sinusoidally-shaped regions of the sinogram associated with pixels using a small MLP. Our results demonstrate that this localized processing framework does significantly improve robustness to OOD data while maintaining nearly constant memory requirements across different resolutions, being computationally efficient even at realistic high resolutions. The differentiable architecture makes it easy to combine Glimpse with other deep learning pipelines. Concretely, we demonstrated the possibility to learn sensor geometry for uncalibrated systems and to adapt the learned filter according to the noise level and projection sparsity.
While the memory required by Glimpse varies little with image resolution, a drawback of the current scheme is that memory and computing costs increase as the number of projections $r$. A possible alternative to the standard MLP architecture which is the culprit for this is to use mixture-of-experts layers [43, 44, 45], which selectively employ smaller MLPs for processing inputs. This mixture-of-experts approach is an effective drop-in replacement for standard MLP layers of language transformers [46] and vision transformers [24].
Glimpse could be integrated with various imaging problems involving line integrals in forward operators such as photoacoustic [47, 48] and cryo-electron tomography (cryoET) [49, 50]. Its future full-3D adaptation may yield efficient architectures which resolve the fundamental memory issues with applications of deep learning in 3D medical imaging. This extension is particularly interesting given the ability of Glimpse to operate locally and its near-fixed memory requirement across resolution, which makes it an ideal choice for large 3D objects.
-A Network Architecture and Training Details
-B Learned Sensor Geometry
CT imaging algorithms such as FBP [14], SART [19], LGS [7], LPD [8] assume that the projection angles $\{\alpha_{m}\}_{m=1}^{r}$ are known. In an uncalibrated system where sensor geometry is different from what the algorithms assume, the quality of reconstruction deteriorates [20, 53]. Glimpse allows directly optimizing the projection angles during training. We thus jointly optimize $\{\alpha_{m}\}_{m=1}^{r}$ with other trainable parameters in (10). This additional angle estimation incurs a very modest computational cost.
In the absence of calibration, we cannot expect to have paired ground truth images. In the following experiments, we only want to showcase the possibility to differentiably optimize over angles in Glimpse so we assume having access to paired data (while simulating the uncalibrated forward operator). In practice we could use a self-supervised loss, for example, based on equivariance [54].
@@ V Limitations and Conclusion
We assess the performance of Glimpse in situations with mismatched projection orientations. In the following experiments, we place $r=30$ sensors uniformly around the object at angles $\alpha=0^{\circ},6^{\circ},...,174^{\circ}$. We conduct a comparative analysis of three models: 1) Glimpse (vanilla), with no learnable sensor geometry, 2) Glimpse (LSG), incorporating the proposed learned sensor geometry, and 3) Glimpse (calibrated), operating under ideal conditions with no model mismatch (informed with correct projection angles). We let the Glimpse (LSG) learn the projection angles from the training data where the optimized values $\{\alpha_{m}\}_{i=1}^{r}$ obtained through training can provide a reliable estimate of the actual projection angles.
Uncalibrated system with random sensor shifts: As shown in Figure 7(a), we randomly perturb projection angles by a normally distributed error so that $\alpha_{i}^{\text{given}}=\mathcal{N}(\alpha_{i},\sigma^{2})$; we set $\sigma=2^{\circ}$. We initialize the projection angles $\{\alpha_{m}\}_{i=1}^{r}$ in the Glimpse (LSG) architecture with $\alpha_{i}^{\text{given}}$. Figure 7(b) shows the estimated projection angles obtained through training—Glimpse (LSG) accurately recovers the angles even in the presence of 30 dB measurement noise. As shown in Figure 7(c), this accurate estimation of projection angles results in high-quality reconstructions by Glimpse (LSG) comparable with the network trained in an ideal calibrated system.
Blind inversion with no information from projection angles: We consider the blind scenario where the model operates without any prior knowledge of the sensor geometry making inversion challenging. As shown in Figure 8(a), we initialize the projection angles $\{\alpha_{m}\}_{i=1}^{r}$ in the Glimpse (LSG) architecture with random values. The estimated projection angles are shown in Figure 8(b), highlighting Glimpse (LSG)’s ability for data-driven sensor geometry estimation. Figure 8(c) presents the reconstructions achieved by Glimpse in both its vanilla and LSG versions. As expected, FBP and the Glimpse (vanilla) show poor reconstructions due to the missing sensor geometry information. On the other hand, Glimpse (LSG) could accurately reconstruct both in-distribution and OOD samples. Remarkably, these results are comparable to those achieved by the calibrated Glimpse with informed projection angles.
@@
-C Optimal Filter for FBP Reconstruction